The gorilla, feeling exuberant, throws a 0.50kg banana straight up in the air from the ground with an initial energy of 30j. How high will the banana travel up?


1
Expert's answer
2022-02-18T15:54:07-0500

The initial kinetic energy is converted into potential energy.

Let's write down the law of conservation of energy.

"E_k=E_p"

"E_k=m\cdot g\cdot h"

Where

"h=\frac{E_k}{m\cdot g}=\frac{30}{0.5\cdot 9.81}=6.12m"